# Hierarchical Prompting Taxonomy: A Universal Evaluation Framework for Large Language Models

### v2

@@ B.4 Level 4: Least-to-Most Prompting
Table 10: Prompt templates of different datasets for Least-to-Most Prompting
B.5 Level 5: Generated Knowledge Prompting
Generated Knowledge prompting represents one of the most complex prompting techniques present in the HPF, where the LLM not only addresses the given task but also incorporates additional relevant information to enhance its response. This method involves prompting another LLM to produce auxiliary knowledge that can be used to better understand and solve the problem. By creating a rich context and leveraging self-generated insights, the LLM can provide more detailed, accurate, and nuanced answers, showcasing the full extent of its capabilities. Table 11 presents the prompt templates utilized for Least-to-Most Prompting across all datasets in our experiments.
@@ B.5 Level 5: Generated Knowledge Prompting
knowledge generatiuon prompt: Generate definitions in french of each word in the text: "english text".
@@
Table 11: Prompt templates of different datasets for Generated Knowledge Prompting
B.6 Adaptive HPF
The prompt-selector in adaptive HPF selects the prompting level based on the task complexity to address the task. Llama 3 8B serves as the prompt-selector in the experiments. The prompt template was meticulously designed to ensure maximum clarity, aiming to reduce hallucinations and select the most effective prompting strategy.
  • Role Prompting: Defines a role for the model in solving the task.
  • Least-to-most prompting: Uses a sequential method to derive essential insights from the task in order to solve it.
Figure 8: HP-Scores across all datasets and LLMs for manual HPF and Adaptive HPF.
Appendix C Human Annotation Policy
To ensure the quality of HP-Scores scored by human experts for the datasets, we implemented a rigorous human annotation process. Human annotators were tasked with calculating the HP-Score for each sample in a given dataset. The HP-Score relative to human experts is a measure of the expert capability in addressing the task based on criteria of HPT. Each sample was scored on a scale from 1 (Lowest prompting level) to 5 (Highest prompting level) for the following criteria:
C.1 Criteria
  1. 1.
    Basic Understanding and Reproduction
    • This criterion evaluates the annotator’s ability to comprehend and accurately reproduce the content. A high score indicates clear and precise reproduction of the information is required for solving the task.
  2. 2.
    Understanding and Interpretation
    • This criterion assesses the depth of understanding and the ability to interpret the information correctly. A high score indicates a detailed and precise understanding of the task is necessary to solve the task.
  3. 3.
    Analysis and Reasoning
    • This criterion measures the ability to analyze the information and apply logical reasoning. A high score demonstrates the need to analyze the information to solve the task.
  4. 4.
    Application of Knowledge and Execution
    • This criterion evaluates the practical application of knowledge and the execution of tasks based on the relevant knowledge. A high score indicates effective application and the requirement of knowledge to solve the task.
C.2 Scoring Process
Each sample of every dataset was scored individually by the annotators on a scale from 1 to 5, where:
  • 1 = Poor
  • 2 = Fair
  • 3 = Good
  • 4 = Very Good
  • 5 = Excellent
The final HP-Score for each dataset was calculated by taking the mean of the values from these four criteria. This approach ensures a comprehensive evaluation of each sample by considering multiple performance criteria.
Appendix D Additional Results
D.1 HP-Scores Interpretation
Figure 8 visualizes HP-Scores for different datasets across various LLMs, comparing manual and adaptive frameworks. Here’s a detailed summary:
For the BoolQ dataset, the manual HPF HP-Scores are generally lower across all LLMs compared to the Adaptive HPF. The adaptive scores are significantly higher, indicating better performance with the manual HPF for solving BoolQ.
For the CSQA dataset, manual HPF HP-Scores are moderate, slightly higher than those for BoolQ. However, similar to BoolQ, the Adaptive HPF HP-Scores are consistently higher, again showing the superiority of the manual HPF over adaptive HPF in solving CSQA.
In comparison to the other datasets, the manual HPF HP-Scores for the IWSLT (0.15) and IWSLT (0.20) datasets both show a significant increase. This increase is particularly noticeable for LLMs such as Phi 3 3.8B and Gemma 7B, which indicate a lack of ability to handle IWSLT. With comparable evaluation scores to manual HPF, which suggests no improvement in adaptive HPF, the adaptive HPF HP-Scores continue to be consistently high.
For both SamSum (0.15) and SamSum (0.20) datasets, the manual HPF HP-Scores are notably high for LLMs such as Phi 3 3.8B and Gemma 7B. Adaptive HPF HP-Scores are consistently high compared to manual HPF HP-Scores.
As evidenced by their lower HP-Scores, manual HPF routinely outperforms adaptive HPF in all tasks and LLMs, including BoolQ, CSQA, and SamSum, where it can outperform human experts. With the exception of solving SamSum (0.15) by Mistral 7B in Adaptive HPF, no LLM was able to complete any task with a HP-Score of less than 5, suggesting the need for additional prompting strategies in HPF or improved framework to avoid hallucinations in selecting prompting level by prompt-selector.
D.2 Hallucination in Adaptive HPF
Hallucinations in prompt-selector refer to instances where the LLM generates incorrect or misleading prompting levels, or nonsensical information that is not supported by HPF. These hallucinations can occur across various tasks, including question answering, multiple-choice questions, translation, and summarization. Below, we discuss the occurrence of hallucinations in the Llama 3 8B which was utilized as prompt-selector in our experiments across different tasks and iterations, highlighting how the LLM’s performance varies and where it tends to hallucinate.
As shown in Figure 9, for the BoolQ task, the prompt-selector initially struggles, indicated by the iterations where it reaches Level 4 with hallucinations (represented by ’…’). However, by the fourth iteration, Llama 3 8B manages to answer correctly at Level 2. For the CSQA task, prompt-selector exhibits hallucinations initially, shown by Level 4 and Level 0 (not included in HPF) responses. Eventually, it corrects itself by the third iteration, providing the correct answer at Level 2. For the IWSLT task, prompt-selector demonstrates a consistent pattern of hallucinations across multiple iterations. Even though Llama 3 8B attempts the translation at Level 2 multiple times, it ultimately fails to provide a correct translation, indicating a persistent hallucination. For the SamSum task, prompt-selector shows initial hallucinations in the first three iterations (Level 4). However, by the fourth and fifth iterations, the prompt-selector starts producing lower levels, finally, Llama 3 8B achieving correct answer at Level 2 in the last iteration .
The results indicate that Llama 3 8B exhibits hallucinations in selecting prompting levels across various tasks and iterations resulting in higher HP-Scores for Adaptive HPF, with performance varying significantly. While the LLM can eventually produce correct answers, as seen in the BoolQ and SamSum tasks, it often requires multiple attempts and may still fail in tasks like IWSLT translation.
Figure 9: Hallucination of prompt-selector in choosing the most effective prompting level to solve corresponding tasks by Llama 3 8B.
Appendix E Additional Experiments
E.1 Increasing Threshold for IWSLT and SamSum Datasets
When comparing the performance of various LLMs, Llama 3-8B stands out for its superior results on the IWSLT and SamSum tasks, consistently performing better in both manual and adaptive frameworks. At a threshold of 0.2, the performance of all LLMs reached saturation for these tasks, indicating that increasing the threshold further did not yield improvements. Both adaptive and manual frameworks struggled with thresholds higher than 0.2, failing to meet or exceed this benchmark for most samples within their iterations. Even the most capable LLM, Llama 3-8B, was unable to address the majority of samples across any prompting level within the HPF. This demonstrates the limitations of current HPFs at higher thresholds and highlights the ongoing challenges in achieving better performance in LLMs.
E.2 Adaptive HPF with Varying Prompt-selectors
Experiments demonstrated that Adaptive HPF is most effective when Llama 3 8B is used as a prompt-selector compared to LLMs, attributed to its ability to produce the lowest number of hallucinations when selecting the prompting level within the iterations of the adaptive HPF. Conversely, Gemma 7B, which exhibited numerous hallucinations across most tasks, was found to be the least effective adaptive HPF. Consequently, the choice of the prompt-selector emerges as a crucial factor influencing the efficiency of the adaptive HPF.

### v4

@@ H.4 Level 4: Least-to-Most Prompting
GSM8k Analyze the question: "question". Break the question into sub-problems: "question". Calculate answers for the subproblems of the question: "pred". Calculate the numerical answer to this question: "question" based on the previous calculations: "pred"
HumanEval Analyze the code: "code". Break the question into sub-problems: "code". Complete code for the subproblems of the code: "pred". Complete the code based on the mentioned constraints: "code" based on the previous calculations: "pred"
MMLU Analyze the question: "question". Elaborate about each option for the question: "question", options: A. "option 1" B. "option 2" C. "option 3" D. "option 4". Based on the analysis: "question", Discard wrong answers among the options: A. "option 1" B. "option 2" C. "option 3" D. "option 4".
Table 12: Prompt templates of different datasets for Least-to-Most Prompting.
H.5 Level 5: Generated Knowledge Prompting
Generated Knowledge prompting is one of the most complex techniques in HPF, where the LLM not only addresses the task but also integrates relevant additional information to enhance its response. This method prompts another LLM to produce auxiliary knowledge, creating a richer context for understanding and solving the problem. By leveraging self-generated insights, the LLM can deliver more detailed, accurate, and nuanced answers. Table 13 shows the prompt templates used for Generated Knowledge Prompting across all datasets in the experiments.
@@ H.5 Level 5: Generated Knowledge Prompting
knowledge generation prompt: Generate definitions in french of each word in the text: "english text".
@@
GSM8k Based on the question:"question", calculate the numerical answer to the question using an interpretation of the question:"pred"
HumanEval Complete the code based on the mentioned constraints:"code" using knowledge of the constraints:"pred"
MMLU Choose the answer. "question", options: A. "option 1" B. "option 2" C. "option 3" D. "option 4" using knowledge of the question:"pred"
@@
Table 13: Prompt templates of different datasets for Generated Knowledge Prompting.